Find sec θ, if the terminal side of the angle passes through the
point (- square root of 5,2)

Answer by lwsshak3(11628)

Find sec θ, if the terminal side of the angle passes through the
point (- square root of 5,2)
***
you are working with a reference right triangle in quadrant II where cos<0.
hypotenuse=√(√5)^2+2^2)=√(5+4)=√9=3
adjacent side=-√5
sec θ=hypotenuse/adjacent side=-3/√5=-(3√5)/5